OpenStudy (anonymous):

Susan begins counting backward from 1298 by 4's, saying one number every 5 seconds. At the same time, Jim begins counting forward from 171 by 3's, saying one number every 5 seconds. What number will they both say at the same time? (A) 640 (B) 644 (C) 648 (D) 650 (E) 654

OpenStudy (perl):

for the first one you can do 1298 - 4 ( t/5) , if you want to use t = seconds

OpenStudy (danjs):

Susan decreases by 4 every 5 seconds t = 5 seconds intervals Susans # = 1928 - 4*t

OpenStudy (danjs):

both count in the same interval of 5 seconds

OpenStudy (anonymous):

So what would the answer be??

OpenStudy (danjs):

Jims # = 171 + 3*t

OpenStudy (danjs):

you want jims # to equal susan's # Jim # = susan #

OpenStudy (perl):

its simpler to use t = 5 second intervals :) so i agree with Dan

OpenStudy (anonymous):

soooooo the answer??

OpenStudy (danjs):

Jim # = Susan # 171 + 3t = 1928 - 4t solve for time t first

OpenStudy (danjs):

I can help you get to the answer, im not paid to do the work for you

OpenStudy (anonymous):

how much do you want?

OpenStudy (anonymous):

lol jk

OpenStudy (danjs):

ahahah

OpenStudy (danjs):

i wish,

OpenStudy (anonymous):

ok thx

OpenStudy (danjs):

can you solve that for t?

OpenStudy (danjs):

171 + 3t = 1928 - 4t

OpenStudy (anonymous):

i believe sooo

OpenStudy (anonymous):

is the answer 648?

OpenStudy (danjs):

i misstyped one number 171 + 3t = 1298 - 4t i get t = 161

OpenStudy (danjs):

so after 161 intervals of counting , each person will be at the same number

OpenStudy (anonymous):

so the answer is 654

OpenStudy (danjs):

From Above: Susans # = 1298 - 4t = 1298 - 4(161) = 654 Jims # = 171 + 3(161) = 654 They both meet at 654 after 161 intervals of counting, or 161*5 = 805 seconds

OpenStudy (anonymous):

OK thanks a lot

OpenStudy (danjs):

welcome

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I became a fan

OpenStudy (danjs):

thanks, feel free bother me whenever i am on with @DanJS